APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Wednesday January 16, 2008 (MJD 54482) ---=== OBSERVING TEAM ===--- Viktor Malanushenko (day) Elena Malanushenko (observer) Audrey Simmons (observer) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- We were not able to observe anything tonight because of snow, clouds, high humidity, and extremely low temperatures (-10.5C). However, we adjusted the fiber and wavelength values in accordance with spectro monthly checkout for sp1 since SOS was reading yellow values in arcs and flats. We attmepted to gather dithered flats several different times, but an error repeatedly came back saying that the mechanicals did not respond correctly. Maybe the motors had trouble because of the cold temperatures? ---=== OBSERVING LOG ===--- Afternoon: ---------- All equipment are checked and work properly. Spectro collimation set for -6.0C. Ventilation is started at 00:12Z. Cart 2 is mounted. Night: ------ 00:35Z - Opened enclosure to very cloudy skies. We attempted to observe through openings in the clouds. 00:57Z - goToField with Cart 2, segue169 plate 2851. We couldnt find any of the guider stars, nor an fk5 star because of the thick clouds. We kept trying. 01:20Z - vs+ voltage watcher error appeared on account of CCD55. 01:32Z - We decided to close because of clouds and nearby radar returns. We left the telescope ventilating. 06:25Z - Stopped ventilating. Radar returns were close by. 06:30Z - Radar sensor went off. It began snowing. Meanwhile, we adjusted the fiber and wavelegnth values for spectro monthly checkout since SOS repeatedly gave yellow values inside flat and arc exposures. We adjusted the fiber values by +3 pixels and moved the wavelength values redder by 3 pixels for sp1. We did not have to adjust sp2 since it was in spec. We ran collimate again and verified that they were set for the night time temperature. We noticed that there is a discrepency with the temperature readings from the specrographs and the temperature inside the enclosure. Right after we closed, we issued a mechSend s and the spectrograph temperatures read -3 and -2 but the enclosure temperature from issuing the "caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA" command read -6 and -7C (note the outside temperature was around -8C). Are the spectrographs reading a wrong value? 08:50Z - Took dithered flats, starting with exposure 49049 and ending at 49058. After the tenth flat, and error message came back: "Finished with specFlat 10 of 15. Setting up Collimator 2008-01-17 09:30:41Z: Original command was !mb100! 2008-01-17 09:30:41Z: expected reply was !mb100 OK! motor b: Mechanicals did not respond correctly; response was mb100 invalid keyed list format: list contains an empty field entry Error: problem with motors (see previous error message) Original motor positions: {a -15424} {b -16010} {c -16610} Current motor positions: {a -15313} {b -15898} {c -16610}" 09:39Z - We reset the motor positions back to their original positions and tried again. sp1>mechMoveColl -311 -312 200 sp2>mechMoveColl -223 -200 210 We began with exposure number 49059 and ended at 49067. It returned the same error message after 8 flats were taken. We returned the collomators again to their original positions but decided not to try again. Maybe the motors are having trouble with the cold temperature? There is no chance of opening. There are radar returns surrounding us still, snow is falling, humidity is still too high, and the temperature is at -10C. We called it a night. 11:27Z - endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7287 22:38Z 23:10Z 100 O 164.40 172.51 51 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): yes QA Procedures Done (y/n): no UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 49027, plate -9999 ------- 22:18 49027 0.0 bias 22:28 49028 0.0 bias ----- sequence 49029, plate 2637 ------- 23:11 49029 10.0 flat 23:14 49030 2.0 arc 23:19 49031 3.0 arc 23:21 49032 2.0 arc 23:58 49033 3.0 arc 00:01 49034 3.0 arc ----- sequence 49035, plate 2851 ------- 00:05 49035 10.0 flat 00:08 49036 2.0 arc ----- sequence 49037, plate -9999 ------- 00:21 49037 0.0 bias 00:24 49038 0.0 bias ----- sequence 49039, plate 2851 ------- 05:21 49039 10.0 flat 05:23 49040 2.0 arc 05:28 49041 3.0 arc 05:30 49042 3.0 arc 06:41 49043 10.0 flat 06:44 49044 2.0 arc 07:11 49045 10.0 flat 07:27 49046 2.0 arc 08:13 49047 3.0 arc 08:16 49048 3.0 arc 08:51 49049 10.0 flat 08:55 49050 10.0 flat 09:00 49051 10.0 flat 09:04 49052 10.0 flat 09:08 49053 10.0 flat 09:13 49054 10.0 flat 09:17 49055 10.0 flat 09:21 49056 10.0 flat 09:25 49057 10.0 flat 09:29 49058 10.0 flat 09:52 49059 10.0 flat 09:56 49060 10.0 flat 10:00 49061 10.0 flat 10:04 49062 10.0 flat 10:08 49063 10.0 flat 10:13 49064 10.0 flat 10:17 49065 10.0 flat 10:21 49066 10.0 flat 10:36 49067 10.0 flat 10:39 49068 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:09Z 26 2 4 286 (WNW) 583 - Cloudy 22:40Z 26 4 12 271 (W) 660 - " 23:10Z 25 3 9 263 (W) 567 - " 23:42Z 25 4 5 264 (W) 603 - " 00:12Z 23 3 8 235 (SW) 593 - " 00:54Z 21 5 6 185 (S) 660 - " 01:26Z 21 5 11 186 (S) 657 - " 01:58Z 20 5 13 145 (SE) 626 - " 02:31Z 17 7 5 126 (SE) 885 - " 03:03Z 15 8 9 121 (ESE) 1033 - " 03:35Z 18 7 10 179 (S) 850 - " 04:07Z 17 7 4 183 (S) 879 - " 04:39Z 17 8 8 179 (S) 992 - " 05:11Z 16 8 8 180 (S) 1004 - " 05:43Z 14 5 8 171 (S) 926 - " 06:15Z 15 5 8 183 (S) 973 - Snowing 06:48Z 15 10 10 176 (S) 4051 - " 07:20Z 14 13 11 202 (SSW) 3781 - " 07:52Z 14 13 15 199 (SSW) 1615 - " 08:24Z 14 13 9 217 (SW) 2428 - " 08:56Z 14 13 5 194 (SSW) 1251 - " 09:28Z 13 12 7 226 (SW) 1733 - " 10:01Z 13 12 6 257 (WSW) 3936 - " 10:31Z 13 11 0 213 (SSW) 4522 - " 11:01Z 14 12 5 262 (W) 1067 - " 11:32Z 14 12 8 264 (W) 1022 - ---=== TELESCOPE STATUS ===--- 00:12Z - Began ventilation. Doors and louvers opened, fans on. 00:35Z - Enclosure off. 00:32Z - Enclosure on due to threat of clouds/snow, but we are still ventilating. 06:25Z - Stopped ventilating. Doors and louvers are closed and fans are off. No interlocks bypassed. Status at: Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 2 Counterweights at: 260 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 254.60 lbs, 19.06 psi Imager 127.60 lbs, 21.47 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_37 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2160 -5.2160 Axial B -0.2142 -0.2142 Axial C -0.0015 -0.0016 Trans D 11.8410 11.8540 Lateral E 10.0711 10.0584 Lateral F 11.6586 11.6586 GALILS Commanded: 4600. -2000. 2650. -12450. -6500. -6550. Actual: 4600. -2000. 2650. -12439. -6477. -6546. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 100.00 -200.00 PriOrient: 0.00 0.00 0.00 99.59 -200.69 SECONDARY: ---------- Focus: 0.00 Air Temp.: -4.0 Alt.: 30.0 MIGS TONIGHT NOMINAL Axial A 0.8033 0.8025 Axial B 0.6131 0.6120 Axial C 0.5487 0.5474 Trans D 0.6005 0.6008 GALILS Commanded: 1627227. 1652165. 1469646. -5900. -7800. Actual: 1627496. 1653168. 1470885. -5887. -7780. SETMIR VALUES SecDesOrient: 1257.00 60.00 -4.00 -20.00 118.90 SecOrient: 1256.33 60.16 -4.26 -20.09 119.17 ---=== PROBLEMS IN DETAIL ===--- Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00049027.fit -5.2 -4.1 -2.7 2.5 6.9 8.8 10.2 sdR-b2-00049027.fit -8.0 -5.7 -4.4 1.8 6.9 9.2 10.7 sdR-r1-00049027.fit -6.3 -5.2 -3.3 1.7 7.7 8.9 11.1 sdR-r2-00049027.fit -7.4 -5.8 -4.3 1.6 7.3 8.9 11.5 ------------------- Initial MechSends: sp1> -19906,-19890,-20992 sp2> -10085, -11042, -11184 T current 0.0C, expected T=-5C Collimate results sp1> mechPiston +6232 steps; sp2> +6763 steps Movements sp1> +1500 and sp2> +1500 New MechSend s: sp1> -18405,-18386,-19491 sp2> -8584, -9532, -9684 ------------